Open Tasks

Tasks currently on our agenda:

  • Remove obsolete cables - still a topic wink (Everybody who is familiar with the old setup)
  • Install temperature sensors for RICH power supplies
  • Re-Install MDC Analog Out connection to Counting House (Erwin)
  • Add fans where missing
  • Install trigger cables TOF-MDC34
  • Install trigger cables for RPC (Alberto)
  • Install trigger cables for Forward Wall
  • Adjust all FanPW to final settings
  • Connect power for RPC switches
  • Install RPC Hadcon (Alberto)
  • MDC - what is missing?

Logbook

18.01.2010
  • Put cables and filters on 10 more new fans (Joern, Jan)
  • Prepare RPC / Shower trigger cables (Gosia)
  • Lay 48V distribution for TRBs MDC12 RICH (Torsten)
  • Lay 48V distribution for TRBs left front (Torsten)
  • Lay 48V distribution for TRBs right front (Torsten)
  • Lay Optical cables left front corner (Jan)
  • Finish TOF TRB stacks (Joern, Jan)
  • Finish MDC1 sector 4 cables (Gosia, Joern)
  • Find and install available network switches (Sergey)

19.01.2010
  • removed power cables for RICH (Joern, Khaled, Jan)
  • Prepare more 50cm power cables (Khaled, Gosia, Joern)
  • Add filters for Fans where missing (Joern)
  • Lay 48V distribution for RPC TRB (Torsten)
  • Lay 48V distribution for RPC (Torsten)
  • Prepared optical cables for MDC 1/2 (Gosia, Joern, Torsten, Jan)
  • Loaded new kernels on many trbs (Sergey)

20.01.2010
  • Made list of available temperature sensors for MDC1/2 (Khaled)
  • Soldered new temperature sensors (Erik)
  • Removed RICH power cable to tent (Erwin)
  • Installed front left and rear right TRB stacks (Andreas)
  • Installed final Ethernet setup for all three corners of the mainframe (Sergey, Jan)
  • Changed stack for CTS (Jan)
  • Continued tests for MDC (Gosia, Joern)
  • Fixed some mistakes on sector 3 plane 3/4 (Gosia, Erwin)

22.01.2010
  • Layed all optical cables for rear right, front right corner and MDC1/2 / RICH to optical switches, Ethernet switches and CTS (Joern, Erwin, Jan)
  • Installed TRB mountings (Andreas)
  • Mounted TRB stacks (Joern, Andreas, Jan)
  • Soldered trigger distributor boards (Gabi)
  • Soldered all temperature sensors for MDC plane 3&4 (Erik, Gosia)
  • Mounted trigger distributor on TRB stacks (Gosia)
  • Removed MDC Analog Out Cables between Patch panels - has to be redone in a better way (Joern, Erwin)
  • Removed LEMO cables on MDC rack

26.01.2010
  • Lay 48V distribution for TRBs right rear (Torsten)
  • Build holding for optical switches (Andreas)
  • Installed and connected optical switches for MDC1/2 & RICH (Jan)
  • Changed RPC stacks on sector 3 & 5 to final setup (Yurek, Jan)
  • Equipped one more RPC stack (Marek, Jan)
  • Prepared power cables for Trigger distributor board (Jan)

27.01.2010
  • Prepared two new trigger cables for RPC (Kathrin)
  • Change power supply for RPC to final setup (Torsten, Jan)

28.01.2010
  • The last three stacks prepared for RPC (Marek, Sergey)

29.01.2010
  • Prepared six TRB stacks for TOF with Power distributor (Erik)
  • Installed two network switches in rear right corner, including cables (Jan)
  • Labeled most of RPC cables individually (Kathrin)
  • Layed trigger cable (17pair) from CTS to both TOF/MDC34 and RPC/Shower / corner of mainframe (Erwin, Jan)
  • Prepared more 50cm 48V cables (Paul)
  • Put RJ45 connectors on PCB for temperature sensors (Paul, Erik)
  • Started removing temporary RPC setup (Jan)

01.02.2010
  • Removed all temporary RPC setups
  • Uninstalled LV power supplies for making changes
  • Put cables for TOF on sector 3
  • Prepared cables for RPC / Shower to be put into new cable chains.

03.02.2010
  • Three stacks of RPC TRBs successfully tested (CTS, old HUB, readout) and ready for mounting (Marek, Sergey)

09.02.2010
  • Installed cable guide on rear corners (Torsten, Andreas, Jan)
  • Continued with TOF cables (Erwin)

18.02.2010
  • Two stacks with TRBs for RPC built and tested (CTS, old HUB, readout) and ready for mounting (Marek, Sergey)

19.02.2010
  • Four TRBs for TOF tested (CTS, old HUB, readout) and ready for mounting (Marek, Sergey)

22.02.2010 (and before)
  • Install new cable guides to RPC/Shower
  • Soldered temperature sensor hub boards
  • Soldered CTS trigger fanout boards
  • Finished testing & labelling half of temperature sensors for MDC 3/4
  • Lay all DAQ cables to each RPC/Shower sector (1x trigger, 6x optical cable, 1x Ethernet, 1x Temp.Sens. 2x CAN)

07.05.2010
  • Installed Forward Wall Hub
  • Changed trigger setup according to Michael's mail

-- JanMichel - 29 Jan 2010
Topic revision: r13 - 07 May 2010, MichaelTraxler
 
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Hades Wiki? Send feedback
Imprint (in German)
Privacy Policy (in German)